OpenStudy (anonymous):

How do you turn 3 1/3% in to a fraction. The answer is 1/30 but I can't figure out how they got the answer.

OpenStudy (anonymous):

3 and 1/3 %. that is 3 + .3333333...repeating so you have 3.333333333 %...to convert from % to decimal, divide by 100, therefore 3.3333% / 100 = 0.03333333...repeating....

OpenStudy (anonymous):

so, this looks similar to 1/3, except there is a 0 before the 3's...so in the same way that multiplying by 10 will "move" the decimal point to the right, this looks like 1/3 has been divided by 10....or multiplied by 1/10

OpenStudy (anonymous):

so \[(1/3) * (1/10)\]

OpenStudy (anonymous):

which is the answer you mentioned

OpenStudy (anonymous):

some fraction, call it x, = 3 and 1/3 %.....therefore, (x * 100) = 3 and 1/3 percent, so x = (3 and 1/3) / 100. 3 and 1/3 = 10/3. so (10/3) / 100 = 10/(3 * 100) = 10/(3*10*10)...the 10 in the numerator and one 10 in the denominator cancel out, leaving 1/(3*10) = 1/30...

OpenStudy (anonymous):

hope this is clear enough

OpenStudy (anonymous):

oh hey dixie

OpenStudy (anonymous):

That makes it very clear. Thank you! We were making a stupid error.

OpenStudy (anonymous):

np, any more questions?

OpenStudy (anonymous):

Not right now. Thanks again : )

OpenStudy (anonymous):

all right, no problem, if you do, just ask another question on the left